OpenStudy (anonymous):

A barrel of tomato sauce has spilled on a tile floor. The sauce flow can be expressed with the function r(t) = 2t, where t represents time in minutes and r represents how far the sauce is spreading. The spilled sauce is creating a circular pattern on the tile. The area of the pattern can be expressed as A(r) = πr2. Part A: Find the area of the circle of spilled sauce as a function of time, or A[r(t)]. Show your work. (6 points) Part B: How large is the area of spilled sauce after 5 minutes? You may use 3.14 to approximate π in this problem. (4 points)

OpenStudy (campbell_st):

so in the area formula replace r with 2t so \[A(r(t)) = \pi (2t)^2\] then simplify it part b substitute t = 5 and calculate the area

OpenStudy (haseeb96):

it is very easy put r(t) =2t in area formula A(r(t)) = pi*r^2 = 3.14 x (2t)^2 now solve it yourself and get your answer

OpenStudy (haseeb96):

for Part B: you will put t=5 in the answer of Part A

OpenStudy (anonymous):

so just to make sure the final answer for part a 12.56t^2

OpenStudy (haseeb96):

yes it is for Part A

OpenStudy (haseeb96):

for part B as i said above you will just put t=5 at the place of t^2 in Part A's answer

OpenStudy (anonymous):

it would end up being 12.56(5)^2 @Haseeb96